# Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability

### v2

@@ V Tine, a Method for Advancing the Solution Distribution Boundary
According to the design of RPM problems, the attributes of an image within the RPM problem are divided into key attributes related to reasoning and non-key attributes unrelated to reasoning. The design purpose of attributes unrelated to reasoning is to interfere with the solver’s reasoning and test its ability to filter attributes. This paper has discovered a point that cannot be neglected: attributes are the only factor that determines the specificity of RPM images. Any two RPM images with the same key attributes have the same identity in the reasoning process, regardless of how different their non-key attributes are. This implies that any sample whose key attributes pass the verification can be considered a correct solution to the RPM problem. For instance, in the RAVEN problem[15], the angle of an entity is considered a non-key attribute that is unrelated to reasoning, while attributes such as color, size, shape, and position are all key attributes. This leads to the situation where a white, medium-sized hexagon positioned in the center retains its reasoning identity regardless of how it is rotated. No matter how it rotates, it will not be transformed from a correct solution to an incorrect one or vice versa.
@@
The design of correct solutions for Bongard-Logo problems is much more complex. The reasoning identity of a Bongard-Logo sample is determined by the context within the specific instance it belongs to. Among the various attributes of a sample, which are key attributes and which are non-key attributes are influenced by the context within the relevant instance. In Bongard-Logo problems, some instances focus on more macroscopic attributes of the image, such as its closure and concavity & convexity; while other instances focus on more detailed attributes, such as the number of entities in the image and the order of their arrangement. This is different from RPM problems, where the key and non-key attributes are consistent across all instances within the RPM.
@@
Given the differences in determining correct solutions, or in identifying key and non-key attributes, between Bongard-Logo and RPM problems, the application of our proposed Funny approach differs when applied to these two types of problems. We will discuss them separately.
VI-A Funny for RPM problems
@@
Given that the key attributes and non-key attributes in the RPM problem are fixed, with key attributes being key across all instances and non-key attributes being unfavored in any instance, this paper proposes that the representations of samples with the same key attributes can be mapped into a single Gaussian distribution. Within this Gaussian distribution, the key attributes of the samples are identical, but the non-key attributes vary. Consequently, the distribution of representations for all samples in the entire RPM problem is formulated as a mixture of Gaussian distributions. By exposing Valen to this mixture of Gaussian distributions instead of individual sample representations, we effectively clarify the distribution of correct solutions for Valen. Our intention is to achieve the transformation of the sample representation space as illustrated in the Figure 15.
@@
Therefore, the purpose of the Funny method is to plan the distribution of samples in the RPM problem as a mixture of Gaussian distributions, with the key attributes serving as the component means.
The design of Valen’s Funny method, applied to tackle the RPM problem, is quite straightforward. The Funny method separately extracts the key attributes and non-key attributes of RPM images, treating the key attributes as the center and the non-key attributes as biases relative to this center. Therefore, we refer to the representation of the key attributes extracted from the RPM image $x_{i}$ as $center_{i}$, and the representation of the non-key attributes as $bias_{i}$. Interestingly, the sum of $center_{i}$ and $bias_{i}$ can be viewed as samples drawn from a Gaussian distribution $N(center_{i},\sigma^{2})$, which represents a distribution with a mean of $center_{i}$ and an unknown variance $\sigma^{2}$. When the $bias_{i}$ follow a standard normal distribution, this Gaussian distribution can be considered as $N(center_{i},1)$, which is a distribution with a mean of $center_{i}$ and a variance of 1. In this scenario, representations of samples with the same key attributes but different non-key attributes only differ in their biases. This implies that the Funny method successfully maps representations of samples with the same key attributes into the same Gaussian distribution.
@@
As can be seen in the figure 16, Funny encodes the image sample $x_{i}$ of the reasoning problem into two sets of homomorphic multi-viewpoint representations, named $center_{i}$ and $bias_{i}$ respectively. The final multi-viewpoint representation $r_{i}$ of the image $x_{i}$ extracted by the Funny representation extractor is obtained by summing $center_{i}$ and $bias_{i}$.
However, the new representation extractor of the Funny method, which encodes images $x_{i}$ into $center_{i}$ and $bias_{i}$, still faces two challenges: 1. How to ensure that the $center_{i}$ stores key attributes while the $bias_{i}$ stores non-key attributes? 2. Given that the union of key and non-key attributes constitutes all attributes, how to guarantee that the combination of $center_{i}$ and $bias_{i}$ can store all attributes of the image $x_{i}$? These two questions pose challenges to both the center and bias individually and collectively.
This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes. This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
Furthermore, to ensure that the combination of $center_{i}$ and $bias_{i}$ can store all attributes within image $x_{i}$, and consequently guarantee that the $bias_{i}$ stores the image attributes omitted by the $center_{i}$, it suffices to maximize the mutual information between combination of $center_{i}$ and $bias_{i}$ and all attributes within image $x_{i}$. The most direct method to maximize the mutual information between image attributes and image representations is to append an additional reconstruction task. Therefore, when the Valen framework incorporates the Funny method, its structure should be as illustrated in the Figure 17.
@@
In the figure 17, we can observe that the representation of key attribute, which is $\{center_{i}|i\in[1,9]\}$, extracted by the Funny extractor for a $3\times 3$ image matrix $\{x_{i}|i\in[1,9]\}$ is superimposed with the sampling results $\epsilon$ from a Gaussian normal distribution $N(0,1)$ and then fed into the regular Valen framework to undertake subsequent reasoning tasks. This is done to clarify with the help of Valen that the representations drawn from the Gaussian distribution $N(center,1)$ have the same reasoning identity.
Furthermore, in the figure 17, we can observe that Funny introduces three additional regression tasks based on mean squared error (MSE) loss to Valen, which are indicated by dashed lines in red, yellow, and green. The green dashed line indicates the regression task between the representation of non-key attributes, which is $\{bias_{i}|i\in[1,9]\}$, and the random sampling results $\epsilon$ from the standard normal distribution $N(0,1)$. It is worth noting that the $\epsilon$ here needs to be resampled and cannot be the one previously superimposed on the $\{center_{i}|i\in[1,9]\}$. The objective of this green task is to enforce the bias to conform to the standard normal distribution $N(0,1)$. The red and yellow dashed lines indicate the two regression tasks performed by the Funny decoder. The task indicated by the red dashed line aims to reconstruct the corresponding RPM image $x_{i}$ from the representation obtained by adding $center_{i}$ and $bias_{i}$. The task indicated by the yellow dashed line aims to reconstruct RPM images from the representation obtained by adding $center_{i}$ and $\epsilon$.
The objectives of these two reconstruction tasks are vastly different. The reconstruction task indicated by the red dashed line is aimed at maximizing the conditional distribution $P(x_{i}|center_{i}+bias_{i})$ during the training process of the Funny decoder. On the other hand, the reconstruction task indicated by the yellow dashed line is intended to ensure that the output of the Funny decoder follows the distribution of RPM images, thereby guaranteeing that the Funny decoder can effectively function as a distribution transformation function.
The objective of the red reconstruction task, which is to maximize the conditional distribution $P(x_{i}|center_{i}+bias_{i})$, is equivalent to maximizing the mutual information between the attributes of $x_{i}$ and $center_{i}+bias_{i}$. This mutual information can be calculated using the following formula:
@@
where $H(x_{i})$ is the entropy of $x_{i}$, and $H(x_{i}|center_{i}+bias_{i})$ is the conditional entropy of $x_{i}$ given $center_{i}+bias_{i}$. This formula quantifies the amount of information shared between $x_{i}$ and the combination of ${center}_{i}$ and ${bias}_{i}$. Additionally, the operations similar to the yellow reconstruction task, which involve constraining the output of the decoder to follow a certain distribution through MSE-based reconstruction tasks, have already been well-established and widely used in VAE[8].
@@
Our design intention for Funny is quite clear. We hope that, with the assistance of Valen, Funny can compel $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes unrelated to reasoning that could potentially interfere with it. By doing so, we aim to model the distribution of multi-viewpoint representations, generated by the $center_{i}+bias_{i}$, as a mixed Gaussian distribution that aids Valen in its reasoning process. The relationship between Funny and Valen is mutually beneficial and symbiotic. Although subsequent experiments demonstrated that Funny is a significantly effective method, given the fact that obtaining disentangled representations under unsupervised learning is not achievable [40], this paper holds a pessimistic attitude towards the extent of Funny’s decoupling of key and non-key attributes.
VI-B Funny not funny
@@
What is of paramount importance is that, a Funny decoder with a conventional structure cannot bear the aforementioned theoretical framework. If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$. Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
@@
Unfortunately and unfunny, there is a gep between theory and practice. This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen. If we choose the normal ViT as the Funny decoder, the introduced regression task will be completed before the Valen’s reasoning task, because the optimization speed of the regression task is faster than that of the reasoning task at this time. We posit that the premature completion of the regression task may result in the $center_{i}$ storing more non-key attribute. Because we have set up the regression task from $center_{i}+\epsilon$ to $x_{i}$. This paper speculates that this might be the issue: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task. The non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen. An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution. Anyway, this paper does not intend to bridge this gap between theory and practice by adjusting the scaling coefficients among the various terms in the loss function.
@@
Fortunately and funny, this paper presents a novel decoder architecture, which we term the “Half-split Decoder” to address this issue. The structure of our designed Half-split Decoder is illustrated in the Figure 18.
@@
The upper half, separated by the blue dashed line in figure 18, represents the representation extractor of Funny, which has been previously shown in Figure 16. The lower half depicts the structure of our designed Half-split Decoder.
In the figure 18, we can observe that our Half-split Decoder is composed of two identical structures, with each structure’s parameters optimized independently without parameter sharing. It is evident that each individual structure resembles an inverted version of a ViT[17], which we will not delve into further in this paper. These two identical structures regress the multi-viewpoint representation $r_{i}$ into two outcomes, $xu_{i}$ and $xd_{i}$, respectively. Here $r_{i}$ can be derived either by summing ${center}_{i}$ with $\epsilon$ or by summing ${center}_{i}$ and $bias_{i}$, corresponding to the two regression requirements of Funny. However, the core of the Half-split Decoder lies not in these structures but in the two distinct activation functions illustrated in the figure 18. We represent these two activation functions as follows:
@@
In the above equation, it is evident that the value range of $f_{1}(xu_{i})$ lies within $(0.5,1)$, while the value range of $f_{2}(xd_{i})$ belongs to $(0,0.5)$. Although the value ranges of both $f_{1}(xu_{i})$ and $f_{2}(xd_{i})$ differ from the value range of the original RPM image $x_{i}$, which is $(0,1)$, we still compute the MSE loss between $f_{1}(xu_{i})$ and $x_{i}$, as well as between $f_{2}(xd_{i})$ and $x_{i}$. The sum of these two errors is used as the new loss function for the two regression tasks associated with the Funny decoder. This loss function can be represented by the following equation:
@@
When both $xu_{i}$ and $xd_{i}$ in the formula are decoded from ${center}_{i}+{bias}_{i}$, the loss function is used to supervise the Mutual Information $\text{MI}(x_{i},{center}_{i}+{bias}_{i})$. Conversely, when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, the loss function is employed to supervise the distribution that the output of the Funny decoder follows.
This Half-split Decoder can ensure that the loss of the regression task introduced by Funny does not decrease to zero in the early stage of the reasoning task of Valen, thereby preventing the multi-view representation from becoming stagnant.
VI-C Funny for Bongard-Logo problem
Funny is inherently a method that analyzes the key degree of sample attributes and modifies the sample representation space accordingly. Since the key degree of an attribute in Bongard-Logo is determined by contextual information, the previous form of Funny is difficult to directly transfer to the Bongard-Logo problem. Therefore, we innovate the Funny method.
This paper argues that we can seek a series of unlearnable Bongard-Logo image sample mapping functions $F(x_{i})$ that do not alter the key degree of image attributes in the current contextual environment. Additionally, these functions do not change the key attributes of the input image but produce varying degrees of alteration to non-key attributes. Clearly, we have not abandoned the motivation of Funny: to make it explicit to Valen that key attributes determine reasoning identity, while non-key attributes do not affect reasoning identity. The previous form of Funny achieved this by mapping the representations of samples with the same key attributes into a Gaussian distribution with a variance of 1.
So, how do we find such functions? Through Human-in-the-Loop (HITL). HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models. By analyzing the three coexisting concepts in Bongard-Logo, we subsequently discovered that some direct data augmentation methods, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$. Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen. It is noteworthy that, unlike in the Bongard-Logo problem, these data augmentation methods may disrupt the key attributes in RPM problem. In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.

### v3

@@ VI-A Key attributes and non-key attributes
According to the design of RPM problems, the attributes of an image are divided into key attributes that are relevant to reasoning and non-key attributes that serve as distractors. The latter are introduced solely to interfere with the solver and to evaluate its ability to filter irrelevant information. It is critical to note that attributes alone determine the identity of an RPM image: any two images that share identical key attributes have the same identity in the reasoning process, regardless of their non-key differences. Consequently, any sample whose key attributes satisfy the governing pattern is deemed a correct solution. For example, in the RAVEN problem[16] the rotation angle of an entity is treated as a non-key attribute, whereas color, size, shape, and position are key; thus, a white, medium-sized hexagon positioned in the center retains its reasoning identity no matter how it is rotated and cannot be transformed by rotation from a correct solution into an incorrect one, or vice versa.
@@
The design of correct solutions for Bongard-Logo problems is much more complex. The reasoning identity of a Bongard-Logo sample is determined by its instance context, which also dictates the key and non-key attributes. Some instances focus on macroscopic attributes such as closure, concavity and convexity, while others concentrate on detailed attributes such as the number of entities and their spatial order. This contrasts with RPM problems, where key and non-key attributes are consistent across all instances.
@@
Due to the differences between Bongard-Logo and RPM problems in determining correct solutions and in distinguishing key from non-key attributes, we apply tailored variants of our Funny approach to the two problem types; each will be discussed separately.
VI-B Funny for RPM problems
@@
Given that the key attributes and non-key attributes in the RPM problem are fixed, with key attributes being key across all instances and non-key attributes being unfavored in any instance, this paper proposes that the representations of samples with the same key attributes can be mapped into a single Gaussian distribution. Within this Gaussian distribution, the key attributes of the samples are identical, but the non-key attributes vary. Consequently, the distribution of representations for all samples in the entire RPM problem is formulated as a mixture of Gaussian distributions [45]. By exposing Valen to this mixture instead of individual sample representations, we effectively clarify the distribution of correct solutions for Valen. Our intention is to achieve the transformation of the sample representation space as illustrated in Figure 15.
@@
Therefore, the Funny method aims to plan the RPM sample distribution as a Gaussian mixture in which the key attributes define the component means.
The design of the Funny method for tackling RPM problems is straightforward: Funny adopts the image attribute extraction scheme illustrated in Figure 16.
Figure 16: The form of image attribute extraction adopted by Funny.
In this figure, the Funny method extracts the key attributes and non-key attributes of RPM images, treating the key attributes as the center and the non-key attributes as biases relative to this center. Accordingly, we refer to the representation of the key attributes extracted from the RPM image $x_{i}$ as $center_{i}$, and that of the non-key attributes as $bias_{i}$. Interestingly, the sum of $center_{i}$ and $bias_{i}$ can be viewed as a sample drawn from the Gaussian distribution $N(center_{i},\sigma^{2})$, which represents a distribution with a mean of $center_{i}$ and an unknown variance $\sigma^{2}$. When the $bias_{i}$ follow a standard normal distribution, this Gaussian distribution can be considered as $N(center_{i},1)$, which is a distribution with a mean of $center_{i}$ and a variance of 1. In this scenario, representations of samples with the same key attributes but different non-key attributes only differ in their biases, implying that the Funny method successfully maps representations of samples with the same key attributes into the same Gaussian distribution.
@@
As can be seen in this figure, Funny encodes the reasoning image $x_{i}$ into two sets of homomorphic multi-viewpoint representations, termed $center_{i}$ and $bias_{i}$ respectively. The final multi-viewpoint representation $r_{i}$ is obtained by element-wise summation of $center_{i}$ and $bias_{i}$.
However, the new representation extractor of the Funny method, which encodes images $x_{i}$ into $center_{i}$ and $bias_{i}$, still faces two challenges: (1) How to ensure that the $center_{i}$ stores key attributes while the $bias_{i}$ captures non-key attributes? (2) Given that the union of key and non-key attributes constitutes all attributes, how to guarantee that the combination of $center_{i}$ and $bias_{i}$ can store all attributes of the image $x_{i}$? These questions pose challenges to both the center and bias individually and collectively.
This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes. Under this constraint, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
Furthermore, we propose that maximizing the mutual information between $center_{i}+bias_{i}$ and the image $x_{i}$ suffices to ensure that this combined representation retains the complete set of attributes within $x_{i}$, thereby guaranteeing that $bias_{i}$ encodes the attributes omitted by $center_{i}$. The most direct method to maximize this mutual information is to add a reconstruction task. Therefore, when the Valen framework incorporates the Funny method, the overall architecture is as shown in Figure 18.
@@
As shown in this figure, the key attribute representations $\{{center}_{i}|i\in[1,9]\}$ extracted from the $\times3$ image matrix $\{x_{i}|i\in[1,9]\}$ are superimposed with $\epsilon\sim N(0,1)$ and then fed into the Valen framework for subsequent reasoning. This process uses the Valen model to confirm that representations sampled from $N(center_{i},1)$ share the same reasoning identity. Moreover, Funny introduces three additional regression tasks based on mean squared error (MSE) loss to Valen; these tasks are indicated by dashed lines in red, yellow, and green. The green dashed line indicates the regression task between the representation of non-key attributes $\{bias_{i}|i\in[1,9]\}$ and fresh random samples $\epsilon\sim N(0,1)$. The objective of this green task is to enforce $bias_{i}$ to follow $N(0,1)$. The red and yellow dashed lines indicate two regression tasks performed by the Funny decoder. The red task reconstructs the original RPM image $x_{i}$ from the combined representation $center_{i}+bias_{i}$, while the yellow task reconstructs $x_{i}$ from $center_{i}+\epsilon$. The objectives of these two reconstruction tasks are vastly different. The red task maximizes the conditional distribution $P(x_{i}|center_{i}+bias_{i})$ during training, which is equivalent to maximizing the mutual information between the attributes of $x_{i}$ and $center_{i}+bias_{i}$. This mutual information can be calculated using the following formula:
@@
where $H(x_{i})$ is the entropy of $x_{i}$, and $H(x_{i}|center_{i}+bias_{i})$ is the conditional entropy of $x_{i}$ given $center_{i}+bias_{i}$. The yellow task forces the output of the Funny decoder to follow the distribution of RPM image pixels, ensuring that the Funny decoder acts as an effective distribution transformation function. Such MSE-based distribution matching is widely adopted in VAEs [10].
@@
Our design intention for Funny is clear: with the assistance of Valen, Funny compels $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes that could cause interference. Subsequently, the distribution of multi-viewpoint representations produced by $center_{i}+bias_{i}$ is planned as a mixture of Gaussians, thereby feeding back to Valen and enhancing its reasoning. The relationship between Funny and Valen is symbiotic. Although subsequent experiments demonstrated that Funny is a significantly effective method, given the fact that obtaining disentangled representations under unsupervised learning is not achievable [42], this paper holds a pessimistic attitude towards the extent of Funny’s decoupling of key and non-key attributes.
VI-C Funny not funny
@@
What is of paramount importance is that, a Funny decoder with a conventional structure cannot bear the aforementioned theoretical framework. If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, and thus hinders satisfactory regression and mutual-information supervision between $center_{i}+bias_{i}$ and $x_{i}$. Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, we obtain excellent regression, yet Valen’s normal reasoning accuracy decreases rather than increases.
@@
Unfortunately and “unfunny”, there remains a gap between theory and practice. We speculate that this may arise from the regression task in Funny interfering with the parameter optimization of the reasoning task in Valen. If we adopt the standard ViT as the Funny decoder, the optimization speed of the regression task would be relatively fast, which could lead to the introduced regression task nearing completion before the reasoning task of Valen. We argue that this premature convergence is the crux: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task; the non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen. An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution. Anyway, this paper does not intend to bridge this gap between theory and practice by adjusting the scaling coefficients among the various terms in the loss function.
@@
Fortunately and funny, this paper presents a novel decoder architecture, which we term the “Half-split Decoder” to address this issue. The structure of the Half-split Decoder is illustrated in the Figure 19.
@@
The upper half, separated by the blue dashed line, shows Funny’s representation extractor, previously illustrated in Figure 17. The lower half depicts the Half-split Decoder.
As shown in Figure 19, the Half-split Decoder is composed of two identical structures, with each structure’s parameters optimized independently without parameter sharing. Each branch resembles an inverted version of a ViT[21], which we will not discuss further. The two branches regress the multi-viewpoint representation $r_{i}$ into two outcomes, $xu_{i}$ and $xd_{i}$, respectively. Here $r_{i}$ can be derived either by summing ${center}_{i}$ with $\epsilon$ or by summing ${center}_{i}$ and $bias_{i}$, corresponding to the two regression targets of Funny. The core of the Half-split Decoder, however, lies not in these branches but in the two distinct activation functions depicted in Figure 19. We express these two activation functions as follows:
@@
Clearly, $f_{1}(xu_{i})$ ranges over $(0.5,1)$, while $f_{2}(xd_{i})$ ranges over $(0,0.5)$. Although both ranges differ from the original RPM image pixel range $(0,1)$, we still compute the MSE loss between $f_{1}(xu_{i})$ and $x_{i}$, as well as between $f_{2}(xd_{i})$ and $x_{i}$. The sum of these two errors is used as the new loss function for the two regression tasks associated with the Funny decoder. This loss can be expressed as:
@@
When both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+{bias}_{i}$, $\ell_{\text{Funny}}$ supervises the Mutual Information $\text{MI}(x_{i},{center}_{i}+{bias}_{i})$, corresponding to Funny’s red regression task; when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, $\ell_{\text{Funny}}$ supervises the distribution that the output of the Funny decoder follows, corresponding to the yellow regression task. The mismatch among the value ranges of $f_{1}(xu_{i})$, $f_{2}(xd_{i})$, and $x_{i}$ ensures that Funny’s regression loss $\ell_{\text{Funny}}$ does not drop to zero during the early stages of Valen’s reasoning, thereby preventing the multi-view representation $r_{i}$ from becoming stagnant.
VI-D Funny for Bongard-Logo problem
Funny is essentially a method that modifies the sample representation space based on the key status of sample attributes. Since this key status is context-dependent in Bongard-Logo, the original Funny cannot be directly transferred to this task. To this end, we argue that identifying a family of unlearnable Bongard-Logo image sample mapping functions $F(x_{i})$ that preserve the key attributes under the current context while introducing controlled perturbations to non-key attributes effectively extends Funny’s methodology to this problem. Clearly, we retain Funny’s original principle: to explicitly signal to Valen that key attributes determine reasoning identity, while non-key attributes do not affect it.
So, how do we find such functions? Through Human-in-the-Loop (HITL)[43]. HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models[43, 44]. By analyzing the three coexisting concepts in Bongard-Logo, we found that some straightforward augmentation techniques, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$. Subsequently, to apply Funny to Valen on Bongard-Logo, we simply need to apply these augmentations randomly to the training samples $x_{i}$ during standard Valen training. It should be noted that, unlike in Bongard-Logo, the same augmentations can corrupt key attributes in the RPM problem; hence the Funny variant tailored for Bongard-Logo is not transferable to RPM, and vice versa.